# Quantum Channel Testing in Average-Case Distance

### v1

@@ 1 Introduction
Fawzi, Flammarion, Garivier and Oufkir [FFGO23] considered the problem of testing whether a given blackbox $\mathcal{M}$ implements a fixed channel $\mathcal{N}$ or is $\varepsilon$-far from $\mathcal{N}$ in the diamond norm. This task is called testing identity to $\mathcal{N}$, and also called channel certification. In the weakest algorithmic model without ancillae or adaptivity, they proved that $d/\varepsilon^{\Theta(1)}$ queries to the blackbox are necessary and sufficient to test identity to a fixed unitary channel. They also showed that $\tilde{\Theta}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{3/2}/\varepsilon^{2}}\right)$ queries are necessary and sufficient to test identity to the completely depolarizing channel, which maps every $d_{\mathrm{in}}$-dimensional input state to the $d_{\mathrm{out}}$-dimensional maximally mixed state.
However, the polynomial dependence on $d_{\mathrm{in}}$ and $d_{\mathrm{out}}$ in the complexity of these channel testers is unsatisfactory. The goal of property testing is to obtain ultra-fast algorithms that only probe a tiny portion of their input. Indeed, a property is said to be “testable” if it can be tested with complexity that depends only on the proximity parameter $\varepsilon$ and not on the size or dimension of the object. Quantum objects are large, as the dimension of the state space of a collection of $n$ quantum systems scales exponentially in $n$, so it is critical to obtain quantum channel testers that (at worst) query the blackbox channel a number of times polylogarithmic in the dimensions of that channel.
The problem here is that diamond distance is a worst-case distance, defined via a maximization over all input states, so two channels can be far apart even if they behave similarly except in the vicinity of a single input state. It is natural that such behavior cannot be distinguished by a tester that does not consider the action of the blackbox channel on a large part of its input domain. In contrast, testers for Boolean functions measure distance by the fraction of the domain on which two functions differ, and this notion of statistical distance inherently captures average-case behavior. Property testing algorithms in general capitalize on local-to-global phenomena that typically arise in such average-case settings.
@@ 1.1 Hardness of channel testing in diamond distance
To make this precise, recall that the trace norm ${\|X\|}_{1}$ of a matrix $X\in\mathbb{C}^{d\times d}$ is given by the sum of its singular values. For two states $\rho,\sigma\in\mathbb{C}^{d\times d}$, the trace distance $\frac{1}{2}{\|\rho-\sigma\|}_{1}$ generalizes the notion of statistical distance between probability distributions. The trace norm for states induces a corresponding trace norm on the space of superoperators, defined by ${\|\mathcal{L}\|}_{1}\coloneqq\displaystyle\max_{{\|X\|}_{1}\leq 1}{\|\mathcal{L}(X)\|}_{1}$ for a superoperator $\mathcal{L}$. The completely bounded trace norm, more commonly known as the diamond norm, is defined similarly but with the maximum taken over all trivial extensions of the superoperator: for $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$,
@@ 1.2 An average-case analogue of the diamond norm
Thus motivated, we now introduce an average-case analogue of the diamond norm. To begin our exposition, note that a natural approach is to replace the maximum in the definition Eq. 1 of the diamond norm with an expectation: for a superoperator $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$, let
${\|\mathcal{L}\|}_{\mathrm{avg}}\coloneqq\displaystyle\operatorname*{\mathbb{E}}_{\boldsymbol{\psi}}{\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{in}}})\cdot\boldsymbol{\psi}\|}_{1},$ (2)
where the expectation is taken over a Haar random [^7] Throughout the paper, we will use boldface font to denote random variables. (pure) state $\boldsymbol{\psi}\in\mathopen{}\mathclose{{}\left(\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}}\right)^{\otimes 2}$. Now, faithfully to the nature of property testing, a large distance between two channels ensures different behavior on many states (unlike in the diamond norm).
However, ${\|\cdot\|}_{\mathrm{avg}}$ has the undesirable feature of being sensitive to the dimension of the ancillary register. In the definition Eq. 1 of the diamond norm, this register may have dimension $d_{\mathrm{in}}$ without loss of generality [Wat18, Theorem 3.46] in the sense that if $X$ ranges over $\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\otimes\mathbb{C}^{d_{\mathrm{anc}}\times d_{\mathrm{anc}}}$ for some $d_{\mathrm{anc}}\geq d_{\mathrm{in}}$, then it holds for all superoperators $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ that
@@
If an analogous statement were to fail to hold for ${\|\cdot\|}_{\mathrm{avg}}$, then it would not be clear why any one value of $d_{\mathrm{anc}}$ should be better motivated than any other. Furthermore it is not immediately clear that the quantity ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}$ is concentrated around its mean for Haar random $\boldsymbol{\psi}\in\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\otimes\mathbb{C}^{d_{\mathrm{anc}}\times d_{\mathrm{anc}}}$, even for $d_{\mathrm{anc}}=d_{\mathrm{in}}$, and this condition is necessary for ${\|\cdot\|}_{\mathrm{avg}}$ to describe the behavior of a channel on “typical” inputs.
@@
Let $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ be a superoperator, and let $\boldsymbol{\psi}\in\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\otimes\mathbb{C}^{d_{\mathrm{anc}}\times d_{\mathrm{anc}}}$ be a Haar random state where $d_{\mathrm{anc}}\geq\Omega(d_{\mathrm{in}})$. Then $\operatorname*{\mathbb{E}}{\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}=\Theta({\|J_{\mathcal{L}}\|}_{1})$, with high probability ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}\leq O({\|J_{\mathcal{L}}\|}_{1})$, and (under a slightly stronger assumption [^8] Specifically, assuming that ${\|\mathcal{L}\|}_{\diamond}\leq o(d_{\mathrm{in}}{\|J_{\mathcal{L}}\|}_{1})$, which holds in almost all cases by Theorem 4.1. Or alternatively, assuming $d_{\mathrm{anc}}\geq\omega(d_{\mathrm{in}})$ rather than just $d_{\mathrm{anc}}\geq\Omega(d_{\mathrm{in}})$.) with high probability ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}\geq\Omega({\|J_{\mathcal{L}}\|}_{1})$, where the asymptotic notation hides universal multiplicative constants.
For sufficiently large values of $d_{\mathrm{anc}}$, a Haar random state will be close to maximally entangled and therefore Theorem 1.3 will follow immediately from the triangle inequality, but the threshold $d_{\mathrm{anc}}\geq\Omega(d_{\mathrm{in}})$ is far too low for such an argument to go through (as we show in Section A.3) and therefore Theorem 1.3 is nontrivial. The lack of explicit averaging in ${\|J_{\mathcal{L}}\|}_{1}$ makes it a more convenient quantity to work with than ${\|\mathcal{L}\|}_{\mathrm{avg}}$, and with ${\|J_{\mathcal{L}}\|}_{1}$ there is no ambiguity regarding the dimension of the ancillary register, so we take ${\|J_{\mathcal{L}}\|}_{1}$ as our definition of the average-case norm despite being less explicitly “average-case” than ${\|\mathcal{L}\|}_{\mathrm{avg}}$:
@@
We define the average-case imitation diamond (ACID) norm of a superoperator $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ to be the quantity
@@
Montanaro and de Wolf [MdW13, Section 5.2] briefly proposed property testing of arbitrary channels in the ACID norm as well, albeit not by this name and without the motivations we give. That the ACID norm is indeed a norm follows from the fact that the trace norm is a norm. The ACID norm is defined similarly to the diamond norm, except that instead of maximizing over all bipartite input states, the input is fixed to the maximally entangled state. However, this does not mean that optimal channel testing in ACID distance is as simple as optimal state testing in trace distance for the corresponding property of the Choi state, as we will see in Section 1.3 when we discuss channel testing in ACID distance and in particular give dimension-independent upper bounds for certain tasks.
In Section 4 we relate the ACID norm to other quantities of interest. For example, we show that it generalizes average-case distances used in property testing of Boolean functions (i.e. statistical distance) and in property testing of unitary transformations [MO10, MdW13, Chen2023Junta]. This further motivates our definition of the ACID norm, especially since most of these generalizations already come with the “right” multiplicative constant (unlike ${\|\cdot\|}_{\mathrm{avg}}$). Additionally, Montanaro and de Wolf [MdW13, Lemma 25] proved that ACID distance is quadratically related to a distance used by Wang [wang2012] in property testing of POVMs. We also compare the ACID norm to the diamond norm and to the “average-case induced trace norm” $\operatorname*{\mathbb{E}}{\|\mathcal{L}(\boldsymbol{\psi})\|}_{1}$ of a superoperator $\mathcal{L}$. The latter quantity is also an “average-case norm”, but it seems to lack most of the other motivations that we give for the ACID norm. Finally, we observe that the ACID norm shares certain convenient mathematical properties with the diamond norm, and discuss the prospect of proving a version of the quantum fault-tolerance theorem with the ACID norm in place of the diamond norm.
@@
Finally, there are also practical motivations for channel testing in the ACID norm. A primary application of channel testing is to determine whether a quantum device built in a laboratory or supplied by a third party actually implements the target channel it was allegedly designed to implement. In some applications the device will always take as input half of a maximally entangled state—examples include nonlocal games [CHSH69], quantum teleportation [Wil19, Sec. 6.2.4], the encoding scheme in superdense coding [Wil19, Sec. 6.2.3], entanglement dilution [Wil19, Sec. 19], and various protocols for quantum communication over a noisy channel [Wil19, Part VI]—and in these cases the ACID norm describes the trace distance between the actual and desired states of the bipartite system arising from the faultiness of the quantum device.
@@ 1.3 Channel certification and tomography in ACID distance
A channel tester is an algorithm that makes queries to a channel $\mathcal{M}$ and tries to correctly output a single bit of classical information about $\mathcal{M}$. We consider three resources which a channel tester may or may not have access to, given the tendency for quantum systems to decohere over time and lose their quantum properties such as entanglement and superposition. First, ancillae: does the tester have access to a system of arbitrarily large dimension, or only to a $d_{\mathrm{in}}$-dimensional system, barely large enough to apply $\mathcal{M}$ to (and which is reset after measuring the output of $\mathcal{M}$)? Second, coherence: if the tester does have ancillae, can it apply $\mathcal{M}$ on different subsystems of an entangled input state and then perform an entangled measurement on the entire output? Or must the tester partition its system as the tensor product of always-unentangled subsystems with only one query to $\mathcal{M}$ made within any given subsystem? And third, adaptivity: can the input to subsequent queries depend on the output of previous queries, or must all queries be made in parallel?
We now present a series of results on channel testing in the ACID norm, which we prove in Section 6 and which we summarize and compare to previous work in Table 1. Consider the task of testing identity to a fixed channel $\mathcal{N}$. Bădescu, O’Donnell and Wright [BOW19] proved that for all states $\sigma\in\mathbb{C}^{d\times d}$, there is an algorithm that performs an entangled measurement on $O(d/\varepsilon^{2})$ copies of an unknown state $\rho\in\mathbb{C}^{d\times d}$ and decides whether $\rho=\sigma$ or ${\|\rho-\sigma\|}_{1}\geq\varepsilon$ with success probability at least 2/3232/32 / 3.999In fact, they proved the stronger statement that given $O(d/\varepsilon^{2})$ copies of two unknown states $\rho$ and $\sigma$, an entangled measurement can decide whether $\rho=\sigma$ or ${\|\rho-\sigma\|}_{1}\geq\varepsilon$ with success probability at least $2/3$. Thus Theorem 1.5 generalizes to testing equality between two unknown channels given query access to both of them. Since ${\|\mathcal{M}-\mathcal{N}\|}_{J}={\|J_{\mathcal{M}}-J_{\mathcal{N}}\|}_{1}$ by definition and since $J_{\mathcal{M}}$ can be constructed using one query to $\mathcal{M}$, the following is immediate by applying the above algorithm with $\sigma=J_{\mathcal{N}}$ and $\rho=J_{\mathcal{M}}$ (and $d=d_{\mathrm{in}}d_{\mathrm{out}}$):
@@
More generally, the query complexity of testing identity to a channel $\mathcal{N}$ in this model is at most the sample complexity of testing identity to $J_{\mathcal{N}}$, which may be $o(d_{\mathrm{in}}d_{\mathrm{out}}/\varepsilon^{2})$ depending on $\mathcal{N}$. However, even with coherence, this blackbox reduction to state certification may be far from optimal for channel certification. For example, consider the channel $\mathcal{N}:\mathbb{C}^{d\times d}\to\mathbb{C}^{1\times 1}$ that traces out its entire input, i.e. $\mathcal{N}(X)=\operatorname*{tr}(X)$. Since $\mathcal{N}$ is the only channel of these dimensions, testing identity to $\mathcal{N}$ trivially requires zero queries, whereas its Choi state is maximally mixed and so the blackbox reduction to state testing would require $\Omega(d/\varepsilon^{2})$ queries [OW15]. The key observation is that regardless of the dimensions of a channel $\mathcal{M}$, the reduced state on the second subsystem of $J_{\mathcal{M}}$ is guaranteed to be maximally mixed, a fact which the blackbox reduction to state certification does not take advantage of. Furthermore, channel certification algorithms may query $\mathcal{M}$ in ways besides constructing $J_{\mathcal{M}}$, analogously to how classical property testing algorithms may be allowed to query a function on explicitly chosen inputs rather than random inputs; we leave it as an open problem whether there exists a property for which an optimal tester must query $\mathcal{M}$ in ways besides constructing its Choi state.
@@
For all fixed channels $\mathcal{N}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ and $\varepsilon>0$, there is an ancilla-free, non-adaptive algorithm that makes $\tilde{O}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{3/2}\big{/}\varepsilon^{2}}\right)$ queries to a channel $\mathcal{M}$, and decides whether $\mathcal{M}=\mathcal{N}$ or ${\|\mathcal{M}-\mathcal{N}\|}_{J}\geq\varepsilon$ with success probability at least $2/3$.
Theorem 1.7 generalizes a result of Fawzi et al. [FFGO23], who proved the same upper bound (without log factors) in the case where $\mathcal{N}$ is the completely depolarizing channel. We also remove the log factors from Theorem 1.6 when $\mathcal{N}$ is the completely depolarizing channel.
We now give our promised dimension-independent upper bounds for testing identity to certain specific types of channels:
Theorem 1.8 (Ancilla-free upper bounds for erasure, unitary, and pure state replacement channels).
@@
For comparison, recall that channel certification in the diamond norm requires $\Omega(\sqrt{d}_{\mathrm{in}}/\varepsilon)$ queries for erasure channels (Theorem 1.1), $d/\varepsilon^{\Theta(1)}$ queries for unitary channels [FFGO23], and $\Omega(d_{\mathrm{in}}/\varepsilon)$ queries for pure state replacement channels (Theorem 1.2). Along the way to proving Theorem 1.8, we also show that for every channel $\mathcal{N}$, testing identity to $\mathcal{I}\otimes\mathcal{N}$ in ACID distance efficiently reduces to testing identity to $\mathcal{N}$ in ACID distance (Theorem 6.9); we consider this observation to be of independent interest as progress toward instance-optimality (see Section 1.4). We also remark that Montanaro and de Wolf [MdW13, Section 5.2.1] gave an $O(1/\varepsilon^{2})$ bound upper bound for testing whether a channel $\mathcal{M}$ satisfies the property of being unitary or is far from that property in ACID distance, by a blackbox reduction to purity testing on $J_{\mathcal{M}}$.

### v2

@@ 1 Introduction
Fawzi, Flammarion, Garivier and Oufkir [FFGO23] considered the problem of testing whether a given blackbox implements a fixed channel $\mathcal{N}$ or is $\varepsilon$-far from $\mathcal{N}$ in the diamond norm. This task is called testing identity to $\mathcal{N}$, and also called channel certification. In the weakest algorithmic model without ancillae or adaptivity, they proved that $d/\varepsilon^{\Theta(1)}$ queries to the blackbox are necessary and sufficient to test identity to a fixed unitary channel. They also showed that $\tilde{\Theta}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{3/2}/\varepsilon^{2}}\right)$ queries are necessary and sufficient to test identity to the completely depolarizing channel, which maps every $d_{\mathrm{in}}$-dimensional input state to the $d_{\mathrm{out}}$-dimensional maximally mixed state.
However, the polynomial dependence on $d_{\mathrm{in}}$ and $d_{\mathrm{out}}$ in the complexity of these channel testers is unsatisfactory. The goal of property testing is to obtain ultra-fast algorithms that only probe a tiny portion of their input. Indeed, a property is said to be “testable” if it can be tested with complexity that depends only on the proximity parameter $\varepsilon$ and not on the size or dimension of the object. Quantum objects are large, as the dimension of the state space of a collection of $n$ quantum systems scales exponentially in $n$, so it is critical to obtain channel testers that (at worst) query the blackbox channel a number of times polylogarithmic in the dimensions of that channel.
The problem here is that diamond distance is a worst-case distance, defined via a maximization over all input states, so two channels can be far apart even if they behave similarly except near a single input state. It is natural that such channels cannot be distinguished by a tester that does not consider the action of the blackbox channel on a large part of its input domain. In contrast, testers for Boolean functions measure distance by the fraction of the domain on which two functions differ, and this notion of statistical distance inherently captures average-case behavior. Property testing algorithms in general capitalize on local-to-global phenomena that typically arise in such average-case settings.
@@ 1.1 Hardness of channel testing in diamond distance
To make this precise, recall that the trace norm ${\|X\|}_{1}$ of a matrix $X$ equals the sum of its singular values. The trace distance $\frac{1}{2}{\|\rho-\sigma\|}_{1}$ between states $\rho$ and $\sigma$ generalizes the notion of statistical distance between probability distributions. The trace norm for matrices induces a corresponding trace norm for superoperators, defined by ${\|\mathcal{L}\|}_{1}\coloneqq\max_{{\|X\|}_{1}\leq 1}{\|\mathcal{L}(X)\|}_{1}$ for a superoperator $\mathcal{L}$. The completely bounded trace norm, more commonly known as the diamond norm, is defined similarly but with the maximum taken over all trivial extensions of the superoperator: for $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$,
@@ 1.2 An average-case analogue of the diamond norm
Thus motivated, we now introduce an average-case analogue of the diamond norm. A natural approach is to replace the maximum in the definition Eq. 1 of the diamond norm with an expectation: for a superoperator $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$, let
${\|\mathcal{L}\|}_{\mathrm{avg}}\coloneqq\displaystyle\operatorname*{\mathbb{E}}_{\boldsymbol{\psi}}{\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{in}}})\cdot\boldsymbol{\psi}\|}_{1},$
where $\boldsymbol{\psi}\in\mathopen{}\mathclose{{}\left(\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}}\right)^{\otimes 2}$ is a Haar random (pure) state. [^7] Throughout the paper, we will use boldface font to denote random variables. However, ${\|\cdot\|}_{\mathrm{avg}}$ has the undesirable feature of being sensitive to the dimension of the ancillary register. In the definition Eq. 1 of the diamond norm, this register may have dimension $d_{\mathrm{in}}$ without loss of generality [Wat18, Theorem 3.46], in the sense that if $X$ ranges over $\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\otimes\mathbb{C}^{d_{\mathrm{anc}}\times d_{\mathrm{anc}}}$ for some $d_{\mathrm{anc}}\geq d_{\mathrm{in}}$ then
@@
If an analogous statement were to fail to hold for ${\|\cdot\|}_{\mathrm{avg}}$, then it would not be clear why any one value of $d_{\mathrm{anc}}$ should be better motivated than any other. It is also not immediately clear that ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}$ is concentrated around its mean, even when $d_{\mathrm{anc}}=d_{\mathrm{in}}$, and this condition is necessary for ${\|\cdot\|}_{\mathrm{avg}}$ to describe the behavior of $\mathcal{L}$ on “typical” inputs (unlike the diamond norm).
@@
Let $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ be a superoperator, let $d_{\mathrm{anc}}\geq\Omega(d_{\mathrm{in}})$, and let $\boldsymbol{\psi}\in\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\otimes\mathbb{C}^{d_{\mathrm{anc}}\times d_{\mathrm{anc}}}$ be a Haar random state. Then $\operatorname*{\mathbb{E}}{\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}=\Theta({\|J_{\mathcal{L}}\|}_{1})$, with high probability ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}\leq O({\|J_{\mathcal{L}}\|}_{1})$, and (under a slightly stronger assumption [^8] Specifically, assuming that ${\|\mathcal{L}\|}_{\diamond}\leq o(d_{\mathrm{in}}{\|J_{\mathcal{L}}\|}_{1})$, which holds in almost all cases by Theorem 4.1. Or alternatively, assuming $d_{\mathrm{anc}}\geq\omega(d_{\mathrm{in}})$ rather than just $d_{\mathrm{anc}}\geq\Omega(d_{\mathrm{in}})$.) with high probability ${\|(\mathcal{L}\otimes\mathcal{I}_{d_{\mathrm{anc}}})\cdot\boldsymbol{\psi}\|}_{1}\geq\Omega({\|J_{\mathcal{L}}\|}_{1})$, where the asymptotic notation hides universal multiplicative constants.
For sufficiently large values of $d_{\mathrm{anc}}$, a Haar random state will be close to maximally entangled and therefore Theorem 1.3 will follow immediately from the triangle inequality, but the threshold $d_{\mathrm{anc}}\geq\Omega(d_{\mathrm{in}})$ is far too low for such an argument to go through (as we show in Section A.3) so Theorem 1.3 is nontrivial. The lack of explicit averaging in ${\|J_{\mathcal{L}}\|}_{1}$ makes it a more convenient quantity to work with than ${\|\mathcal{L}\|}_{\mathrm{avg}}$, and with ${\|J_{\mathcal{L}}\|}_{1}$ there is no ambiguity regarding the dimension of the ancillary register, so we take ${\|J_{\mathcal{L}}\|}_{1}$ as our definition of the average-case norm:
@@
The average-case imitation diamond (ACID) norm of a superoperator $\mathcal{L}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ is the quantity
@@
Montanaro and de Wolf [MdW13, Section 5.2] briefly proposed property testing of arbitrary channels in the ACID norm as well, albeit not by this name and without the motivations we give. That the ACID norm is indeed a norm follows from the fact that the trace norm is a norm. The ACID norm is defined similarly to the diamond norm, except that instead of maximizing over all bipartite input states, the input is fixed to the maximally entangled state. However this does not mean that optimal channel testing in ACID distance is as simple as optimal state testing in trace distance for the corresponding property of the Choi state, as we will see in Section 1.3.
In Section 4 we relate the ACID norm to other quantities of interest. We show that it generalizes average-case distances used in property testing of Boolean functions (i.e. statistical distance) and in property testing of unitary transformations [Low2009Avg, MO10, Wang2011Unitary, MdW13, Chen2023Junta, Zhao+23]. This further motivates our definition of the ACID norm, especially since the ACID norm already has the “right” multiplicative constant for some of these generalizations (unlike ${\|\cdot\|}_{\mathrm{avg}}$). Additionally, Montanaro and de Wolf [MdW13, Lemma 25] proved that ACID distance is quadratically related to a distance used by Wang [wang2012] in POVM testing. We also compare the ACID norm to the diamond norm and to the “average-case induced trace norm” $\operatorname*{\mathbb{E}}{\|\mathcal{L}(\boldsymbol{\psi})\|}_{1}$ of a superoperator $\mathcal{L}$. The latter quantity is also an “average-case norm”, but it seems to lack most of the other motivations that we give for the ACID norm. Finally, we observe that the ACID norm shares certain convenient mathematical properties with the diamond norm, and discuss the prospect of proving a version of the quantum fault-tolerance theorem with the ACID norm in place of the diamond norm.
@@
Finally, there are also practical motivations for channel testing in the ACID norm. A primary application of channel testing is to determine whether a quantum device built in a laboratory or supplied by a third party actually implements the target channel it was allegedly designed to implement. In some applications the device will always take as input half of a maximally entangled state—examples include nonlocal games [CHSH69], quantum teleportation [Wil19, Sec. 6.2.4], the encoding scheme in superdense coding [Wil19, Sec. 6.2.3], entanglement dilution [Wil19, Sec. 19], and various protocols for quantum communication over a noisy channel [Wil19, Part VI]—and in these cases ACID distance describes the trace distance between the actual and desired states of the bipartite system arising from the faultiness of the quantum device.
@@ 1.3 Channel certification and tomography in ACID distance
A channel tester is an algorithm that makes queries to a channel $\mathcal{M}$ and tries to decide whether $\mathcal{M}$ satisfies or is far from some property. We consider three resources which a channel tester may or may not have access to, given the tendency for quantum systems to decohere over time and lose their quantum properties such as entanglement and superposition. First, ancillae: does the tester have access to a system of arbitrarily large dimension, or only to a $d_{\mathrm{in}}$-dimensional system, barely large enough to apply $\mathcal{M}$ to (and which is reset after measuring the output of $\mathcal{M}$)? Second, coherence: if the tester does have ancillae, can it apply $\mathcal{M}$ on different subsystems of an entangled input state and then perform an entangled measurement on the entire output? Or must the tester partition its system as the tensor product of always-unentangled subsystems with only one query to $\mathcal{M}$ made within any given subsystem? And third, adaptivity: can the input to subsequent queries depend on the output of previous queries, or must all queries be made in parallel?
We now present a series of results on channel testing in ACID distance, which we prove in Section 6 and which we summarize and compare to previous work in Table 1. Consider the task of testing identity to a fixed channel $\mathcal{N}$. Bădescu, O’Donnell and Wright [BOW19] proved that for all states $\sigma\in\mathbb{C}^{d\times d}$, there is an algorithm that performs an entangled measurement on $O(d/\varepsilon^{2})$ copies of an unknown state $\rho\in\mathbb{C}^{d\times d}$ and decides whether $\rho=\sigma$ or ${\|\rho-\sigma\|}_{1}\geq\varepsilon$ with success probability at least 2/3232/32 / 3.999In fact, they proved the stronger statement that given $O(d/\varepsilon^{2})$ copies of two unknown states $\rho$ and $\sigma$, an entangled measurement can decide whether $\rho=\sigma$ or ${\|\rho-\sigma\|}_{1}\geq\varepsilon$ with success probability at least $2/3$. Thus Theorem 1.5 generalizes to testing equality between two unknown channels given query access to both of them. Since ${\|\mathcal{M}-\mathcal{N}\|}_{J}={\|J_{\mathcal{M}}-J_{\mathcal{N}}\|}_{1}$ by definition and since $J_{\mathcal{M}}$ can be constructed using one query to $\mathcal{M}$, the following is immediate by applying the above algorithm with $\sigma=J_{\mathcal{N}}$ and $\rho=J_{\mathcal{M}}$ (and $d=d_{\mathrm{in}}d_{\mathrm{out}}$):
@@
More generally, the query complexity of testing identity to a channel $\mathcal{N}$ in this model is at most the sample complexity of testing identity to $J_{\mathcal{N}}$, which may be $o(d_{\mathrm{in}}d_{\mathrm{out}}/\varepsilon^{2})$ depending on $\mathcal{N}$. However, even with coherence, this blackbox reduction to state certification may be far from optimal for channel certification. For example, consider the channel $\mathcal{N}:\mathbb{C}^{d\times d}\to\mathbb{C}^{1\times 1}$ that traces out its entire input, i.e. $\mathcal{N}(X)=\operatorname*{tr}(X)$. Since $\mathcal{N}$ is the only channel of these dimensions, testing identity to $\mathcal{N}$ trivially requires zero queries, whereas its Choi state is maximally mixed and so the blackbox reduction to state testing would require $\Omega(d/\varepsilon^{2})$ queries [OW15]. The key observation is that regardless of the dimensions of a channel $\mathcal{M}$, the reduced state on the second subsystem of $J_{\mathcal{M}}$ is guaranteed to be maximally mixed, a fact which the blackbox reduction to state certification does not take advantage of. Furthermore, channel certification algorithms may query $\mathcal{M}$ in ways besides constructing $J_{\mathcal{M}}$, analogously to how classical property testing algorithms may be allowed to query a function on explicitly chosen inputs rather than random inputs; we leave it as an open problem whether there exists a channel $\mathcal{N}$ for which an optimal certification algorithm must query $\mathcal{M}$ in ways besides constructing its Choi state.
@@
For all fixed channels $\mathcal{N}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$ and $\varepsilon>0$, there is an ancilla-free, non-adaptive algorithm that makes $\tilde{O}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{3/2}\big{/}\varepsilon^{2}}\right)$ queries to a channel $\mathcal{M}$, and decides whether $\mathcal{M}=\mathcal{N}$ or ${\|\mathcal{M}-\mathcal{N}\|}_{\diamond}\geq\varepsilon$ with success probability at least $2/3$.
Theorem 1.7 generalizes a result of Fawzi et al. [FFGO23], who proved the same upper bound without log factors in the case where $\mathcal{N}$ is the completely depolarizing channel. We also remove the log factors from Theorem 1.6 when $\mathcal{N}$ is the completely depolarizing channel.
We also give dimension-independent upper bounds for testing identity to certain channels:
Theorem 1.8 (Erasure, unitary, and pure state replacement channel certification).
@@
For comparison, recall that channel certification in diamond distance requires $\Omega(\sqrt{d}_{\mathrm{in}}/\varepsilon)$ queries for erasure channels (Theorem 1.1), $d/\varepsilon^{\Theta(1)}$ queries for unitary channels [FFGO23], and $\Omega(d_{\mathrm{in}}/\varepsilon)$ queries for pure state replacement channels (Theorem 1.2). Along the way to proving Theorem 1.8, we also show that for every channel $\mathcal{N}$, testing identity to $\mathcal{I}\otimes\mathcal{N}$ in ACID distance efficiently reduces to testing identity to $\mathcal{N}$ in ACID distance (Theorem 6.7); we consider this observation to be of independent interest as progress toward instance optimality (see Section 1.4). We also remark that Montanaro and de Wolf [MdW13, Section 5.2.1] gave an $O(1/\varepsilon^{2})$ bound upper bound for testing whether a channel $\mathcal{M}$ satisfies the property of being unitary or is far from that property in ACID distance, by a blackbox reduction to purity testing on $J_{\mathcal{M}}$.